# AdaCQR: Enhancing Query Reformulation for Conversational Search via Sparse and Dense Retrieval Alignment

### v1

@@ 3.5.1 Training Stage 1 for Initialization
In the first training stage, we train a language model using the superior reformulation labels to endow it basic capability of query reformulation. To encourage more diverse generation results, a label smooth cross-entropy loss is used:
@@
where $\mathcal{P}$ is the reformulation problem including current query $q$ and historical context $\mathrm{H}$, $Q^{\star}_{<j}$ is the first $j$ tokens of the reformulation label $Q^{\star}$. $p_{s}$ is a label smooth distribution, defined as follows:
@@ 4 Experiments
We train and evaluate our model using two widely utilized conversational search datasets: QReCC  Anantha et al. (2021) and TopiOCQA  Adlakha et al. (2022). The details of these datasets are shown in Appendix B.1.
@@
Following prior works in the CQR task  Wu et al. (2022); Mo et al. (2023a); Jang et al. (2023); Yoon et al. (2024), we evaluate AdaCQR using sparse and dense retrieval systems [^2] We do not fine-tune the retrievers within our framework, thus excluding consideration of such CDR work in baselines.. The sparse retrieval system used is BM25  Robertson et al. (2009). For dense retrieval, we use ANCE  Xiong et al. (2020), trained on the MS MARCO  Nguyen et al. (2016) retrieval task.
TopiOCQA QReCC
Type Query Reform. MRR NDCG R@10 R@100 MRR NDCG R@10 R@100
Sparse (BM25) Human Rewrite - - - - 39.8 36.3 62.7 98.5
T5QR (T5-base) 11.3 9.8 22.1 44.7 33.4 30.2 53.8 86.1
CONQRR (T5-base) - - - - 38.3 - 60.1 88.9
IterCQR (T5-base) 16.5 14.9 29.3 54.1 46.7 44.1 64.4 85.5
InfoCQR (ChatGPT) - - - - 49.4 - 67.1 88.2
ConvGQR (T5-base) 12.4 10.7 23.8 45.6 44.1 41.0 64.4 88.0
RetPO (LLaMA2-7B) 28.3 26.5 48.3 73.1 50.0 47.3 69.5 89.5
AdaCQR (Ours, T5-base) 17.8 15.8 34.1 62.1 52.4 49.9 70.9 91.0
  +Expansion 28.3 26.5 48.9 71.2 55.1 52.5 76.5 93.7
Dense (ANCE) Human Rewrite - - - - 38.4 35.6 58.6 78.1
T5QR (T5-base) 23.0 22.2 37.6 54.4 34.5 31.8 53.1 72.8
CONQRR (T5-base) - - - - 41.8 - 65.1 84.7
IterCQR (T5-base) 26.3 25.1 42.6 62.0 42.9 40.2 65.5 84.1
ConvGQR (T5-base) 25.6 24.3 41.8 58.8 42.0 39.1 63.5 81.8
RetPO (LLaMA2-7B) 30.0 28.9 49.6 68.7 44.0 41.1 66.7 84.6
AdaCQR (Ours, T5-base) 32.8 31.5 54.6 73.0 45.1 42.4 66.3 83.4
  +Expansion 38.5 37.6 58.4 75.0 45.8 42.9 67.3 83.8
[…]
@@
In our study, we compare AdaCQR with the following representative baselines in the CQR task: (1) T5QR  Lin et al. (2020b) is a vanilla baseline that utilizes the T5-base Raffel et al. (2020) model to perform CQR tasks. (2) CONQRR  Wu et al. (2022) aligns the reformulation model with retrievers through direct optimization using reinforcement learning. (3) ConvGQR  Mo et al. (2023a) enhances retrieval performance by employing two generative models, one for query reformulation and the other for query expansion. (4) InfoCQR  Ye et al. (2023) employs ChatGPT  OpenAI (2022) to conduct query reformulation via a “rewrite-then-edit” process. (5) IterCQR  Jang et al. (2023) achieves alignment by minimizing Bayes Risk based on the semantic similarity between the query and the gold passage. (6) RetPO Yoon et al. (2024) utilizes large language models to generate multiple reformulations through multi-perspective prompting, creates binarized comparisons based on retriever feedback, and optimizes LLaMA2-7B Touvron et al. (2023) using direct preference optimization (DPO) Rafailov et al. (2023).
The details regarding implementation and evaluation metrics are provided in the Appendix C and Appendix B.2, respectively.
@@ 4.1 Main Results
To evaluate the efficacy of our framework, we conducted comprehensive experiments on the QReCC and TopicOCQA datasets, employing the AdaCQR model trained individually on each dataset, presented in Table 1. We consider three kinds of backbones as baselines: the T5-based, the LLaMA2-7B-based, and the ChatGPT-based. The results demonstrate that AdaCQR significantly outperforms previous models utilizing T5-base as the backbone. Furthermore, AdaCQR exceeds the performance of RetPO, which uses LLaMA2-7B as the backbone, in both the QReCC and the dense retrieval section of TopiOCQA, underscoring the effectiveness of our approach. Additionally, the enhanced performance of RetPO is attributed to the inherently strong common-sense reasoning capabilities of the backbone model. To a fair comparison with RetPO, we employed the vanilla LLaMA2-7B, generating potential answers or keywords as query expansions for reformulation queries. AdaCQR with query expansion achieves results comparable to RetPO in sparse retrieval on the TopiOCQA, while significantly outperforming RetPO in other settings. Specifically, AdaCQR with expansion shows superior performance in dense retrieval on the TopiOCQA, attaining the best MRR (38.5), NDCG (37.6), R@10 (58.4), and R@100 (75.0).
@@ 4.2 Ablation Study
QReCC
Type Query Reform. MRR R@10
Sparse Superior labels $Q^{\star}$ 45.4 65.5
AdaCQR (Ours) 52.4 70.9
@@
   w/o. Labels $Q^{\star}$ 44.9 63.7
Dense Superior labels $Q^{\star}$ 40.1 60.2
AdaCQR (Ours) 45.1 66.3
@@
   w/o. Labels $Q^{\star}$ 41.0 60.5
@@
Table 2: Ablation study for each component of AdaCQR. We also report the performance of the superior labels $Q^{\star}$ which are obtained by prompting LLMs through in-context learning, as detailed in Section 3.4.
In this study, we employ a contrastive loss to align the retrievers. We also introduce a fusion metric to evaluate query performance across semantic and term perspectives. Additionally, we leverage LLM to obtain superior labels $Q^{\star}$ through prompting, reducing reliance on sub-optimal and costly human labels. To investigate the impact of each component on the performance of AdaCQR, we conducted ablation experiments focusing on these three specific modules in Table 2. To assess the influence of contrastive loss, we executed a single-stage training process without alignment. To determine the effect of the fusion metric, we substituted it with the evaluation method used in previous work Jang et al. (2023), which only relies on the cosine similarity between the query and the gold passage. To further investigate the effectiveness of considering both perspectives in the fusion metric, we separately remove sparse ranking $r_{s}$ and dense ranking $r_{d}$ within it for analysis. To examine the impact of superior labels $Q^{\star}$, we trained AdaCQR using human rewrite labels instead. The results of these ablation experiments reveal that the exclusion of any of these modules greatly affects the performance of AdaCQR, showing the importance of these components for AdaCQR. In particular, the most notable decline in performance occurs upon the removal of Contrastive Loss. Its decline in performance is followed by the impact of the labels $Q^{\star}$ and the Fusion Metric. Removing any rank degrades performance for both retrievers, more significantly for the corresponding retriever. This confirms the rationale behind considering both perspectives simultaneously. It is worth noting that superior labels $Q^{\star}$ can achieve comparable performance both in sparse and dense retrievals, which validates the effectiveness of the proposed fusion metric and the annotation method. The results indicate that queries reformulated by AdaCQR significantly outperform superior labels $Q^{\star}$, demonstrating the advantages of an aligned model in CQR.
4.3 Robustness to Topic Shifts in Conversation
Topic-Concentrated Topic-Shifted
Model MRR R@10 MRR R@10
T5QR 35.2 54.4 25.2 45.1
CONQRR 41.9 63.1 25.2 45.9
IterCQR 54.4 72.4 24.9 49.7
Human Rewrite 44.0 66.7 31.8 56.7
AdaCQR 66.0 82.4 34.1 58.3
Table 3: Performance of AdaCQR on topic-concentrated and topic-shifted samples on QReCC, MRR and R@10 are reported. The result is reported on BM25 Retrieval System.
@@ 5 Analysis
5.1 Effect of the Multi-Task Loss
QReCC
Coefficient($\gamma$) MRR NDCG R@10 R@100
0 43.3 41.0 62.8 88.5
0.1 45.3 42.7 65.2 90.2
1 48.8 46.1 68.7 91.2
10 50.2 47.7 68.8 89.0
100 52.4 49.9 70.9 91.0
1000 49.4 46.7 68.6 90.7
$\infty$ 44.5 41.8 65.5 90.9
@@ 5.2 Analysis of the Aligned Query
Figure 3: Analysis of the aligned reformulation query across different epochs in Stage 2 training, focusing on the term overlap with the gold passage (DICE coefficient), and semantic similarity to the gold passage (cosine similarity).
To evaluate the effectiveness of the aligned reformulation queries, we analyzed the reformulation queries across the first 5 epochs during Stage 2 training in Figure 3. We conducted analyses focusing on the average term overlap and semantic similarity between the queries and the gold passages. The DICE Coefficient Dice (1945) is utilized to assess term overlap, while cosine similarity is employed to measure semantic similarity. This analysis indicates that both term overlap and semantic similarity between the reformulated queries and the gold passages exhibit an increasing trend with each epoch in Stage 2, demonstrating the effectiveness of our method in considering both perspectives.
@@ 6 Conclusion
In this paper, to achieve alignment between the reformulation model and both term-based and semantic-based retrieval systems, AdaCQR is proposed to enhance the generalizability of information-seeking queries across diverse retrieval environments. We developed effective techniques to acquire superior reformulation labels and generate diverse input candidates, boosting the efficiency and robustness of the AdaCQR framework. Extensive experiments on two datasets demonstrate the superiority of AdaCQR, achieving performance comparable with the LLaMA2-7B model while using only the T5-base.
@@
During the AdaCQR training process, we leverage ChatGPT for superior reformulation label annotation, and our annotation prompt requires training a basic model, which incurs additional costs and training expenses. Furthermore, due to budget constraints, we did not use more powerful LLMs, such as GPT-4 to obtain reformulation labels, although it is obvious that employing a more powerful LLM would yield better reformulation labels.
Although no further costs are introduced during reformulation model inference, aligning AdaCQR with retrievers introduces additional training time. Furthermore, generating the sorted candidate set for alignment demands extra retrieval time and increased storage capacity.

### v3

@@ 3.5.1 Training Stage 1 for Initialization
In the first training stage, we train a language model using the superior reformulation labels to endow it with the basic capability of query reformulation. To encourage more diverse generation results, a label smooth cross-entropy loss is used:
@@
where $\mathcal{P}$ is the reformulation session including current query $q$ and historical context $\mathrm{H}$, $Q^{\star}_{<j}$ is the first $j$ tokens of the reformulation label $Q^{\star}$. $p_{s}$ is a label smooth distribution, defined as follows:
@@ 3.5.2 Candidates Generation for Alignment
TopiOCQA QReCC
Type Query Reform. MRR NDCG R@10 R@100 MRR NDCG R@10 R@100
Sparse (BM25) T5QR (T5-base) 11.3 9.8 22.1 44.7 33.4 30.2 53.8 86.1
CONQRR (T5-base) - - - - 38.3 - 60.1 88.9
EDIRCS (T5-base) - - - - 41.2 - 62.7 90.2
IterCQR (T5-base) 16.5 14.9 29.3 54.1 46.7 44.1 64.4 85.5
LLM-Aided (ChatGPT) - - - - 49.4 46.5 67.1 88.2
LLM4CS-REW(ChatGPT) 16.8 15.0 31.3 56.7 35.2 32.0 55.9 84.7
AdaCQR (Ours, T5-base) 17.8 15.8 34.1 62.1 52.4 49.9 70.9 91.0
ConvGQR (T5-base) 12.4 10.7 23.8 45.6 44.1 41.0 64.4 88.0
RetPO (LLaMA2-7B) 28.3 26.5 48.3 73.1 50.0 47.3 69.5 89.5
LLM4CS-RAR(ChatGPT) 27.9 26.4 48.4 71.1 51.6 49.3 75.3 92.6
AdaCQR+Expansion 28.3 26.5 48.9 71.2 55.1 52.5 76.5 93.7
Dense (ANCE) T5QR (T5-base) 23.0 22.2 37.6 54.4 34.5 31.8 53.1 72.8
CONQRR (T5-base) - - - - 41.8 - 65.1 84.7
EDIRCS (T5-base) - - - - 42.1 - 65.6 85.3
IterCQR (T5-base) 26.3 25.1 42.6 62.0 42.9 40.2 65.5 84.1
LLM-Aided (ChatGPT) - - - - 43.5 41.3 65.6 82.3
LLM4CS-REW (ChatGPT) 30.3 29.0 49.9 67.7 36.8 34.0 56.1 74.2
AdaCQR (Ours, T5-base) 32.8 31.5 54.6 73.0 45.1 42.4 66.3 83.4
ConvGQR (T5-base) 25.6 24.3 41.8 58.8 42.0 39.1 63.5 81.8
RetPO (LLaMA2-7B) 30.0 28.9 49.6 68.7 44.0 41.1 66.7 84.6
LLM4CS-RAR (ChatGPT) 35.4 34.4 55.2 72.2 44.7 41.8 67.2 84.0
AdaCQR+Expansion 38.5 37.6 58.4 75.0 45.8 42.9 67.3 83.8
Table 1: Evaluation results of various retrieval system types on the test sets of QReCC and TopiOCQA. The best results among all methods with similar settings are bolded, and the second-best results are underlined. $\heartsuit$ denotes the method involved in using query expansion. $\ddagger$ denotes the baselines utilizing another dual encoder dense retrieval. +Expansion denotes the addition of query expansion, details in Appendix F.
@@ 4 Experiments
We train and evaluate our model using two widely utilized conversational search datasets: QReCC  Anantha et al. (2021) and TopiOCQA  Adlakha et al. (2022). We also conduct zero-shot experiments on TREC CAsT 19-21 Dalton et al. (2020, 2021, 2022). The details of these datasets are shown in Appendix B.2.
@@
Following prior works in the CQR task  Wu et al. (2022); Mo et al. (2023a); Jang et al. (2023); Yoon et al. (2024), we evaluate AdaCQR using sparse and dense retrieval systems The sparse retrieval system used is BM25  Robertson et al. (2009). For dense retrieval, we use ANCE  Xiong et al. (2020), trained on the MS MARCO  Nguyen et al. (2016) retrieval task.
CAsT-19 CAsT-20 CAsT-21
Query Reform. MRR R@10 R@100 MRR R@10 R@100 MRR R@10 R@100
T5QR(T5-base) 70.1 - 33.2 42.3 - 35.3 46.9 - 40.8
EdiRCS(T5-base) 70.9 - 35.3 43.8 - 37.5 - - -
LLM4CS-REW(ChatGPT) 65.6 10.5 33.4 49.1 16.9 41.3 57.7 21.9 55.3
AdaCQR(Ours, T5-base) 71.6 12.1 36.7 51.4 15.7 40.9 58.3 21.3 54.5
ConvGQR(T5-base) 70.8 - 33.6 46.5 - 36.8 43.3 - 33.0
LLM4CS-RAR(ChatGPT) 70.6 12.6 37.7 57.3 19.8 47.6 65.8 24.6 59.5
AdaCQR+Expansion 74.5 13.8 39.2 56.6 19.2 45.6 64.2 25.0 58.7
Table 2: Zero-shot experiment results on TREC CAsT 19-21 datasets. The best results among all methods with similar settings are bolded, and the second-best results are underlined.
@@
To compare with previous CQR baselines, we define two variants of our framework:
  • AdaCQR generates the reformulation queries through the aligned T5-base model, as illustrated in Section 3.5.
  • AdaCQR+Expansion concats the reformulation queries generated by AdaCQR and query expansions generated by vanilla LLaMA2-7B leveraging the pseudo answers and keywords expansion techniques as described in the Appendix F.
We categorize previous works into two groups: query reformulation and query reformulation with expansion. In the query reformulation section, we compare AdaCQR with the fine-tuned T5-base models including: T5QR Lin et al. (2020), CONQRR Wu et al. (2022), EDIRCS Mao et al. (2023a), IterCQR Jang et al. (2023) and LLM-based prompting methods including LLM-Aided Ye et al. (2023) and LLM4CS Mao et al. (2023b) under the Rewriting Prompting(REW) setting. In the query reformulation with expansion section, we compare AdaCQR+Expansion with T5-based model ConvGQR222 AdaCQR (T5-base) without expansion, is showing superior performance than ConvGQR whose expansions are generated by another T5-base model. Mo et al. (2023a), the fine-tuned LLaMA2-7B model RetPO Yoon et al. (2024) and LLM-based prompting method LLM4CS Mao et al. (2023b) with the Rewrite-and-Resonse(RAR) setting.
The details regarding baselines, implementation and evaluation metrics are provided in Appendix B.1, Appendix C and Appendix B.3, respectively.
@@ 4.1 Main Results
To evaluate the efficacy of our framework, we conducted comprehensive experiments datasets with the AdaCQR and baselines, presented in Table 1. We consider three kinds of backbones as baselines: the T5-based, the LLaMA2-7B-based, and the ChatGPT-based.
The results under the without expansion setting demonstrate that AdaCQR significantly outperforms previous models utilizing T5-base as the backbone. Among methods with expansion, the exceptional performance of RetPO and LLM4CS-RAR, which use LLaMA2-7B and ChatGPT as the backbone, respectively, can be attributed to the inherently strong common-sense reasoning capabilities of backbone models. AdaCQR with expansion achieves results comparable to RetPO and LLM4CS-RAR, with only a slight disadvantage in the R@100 metric, while empirically outperforming them in other settings. Specifically, AdaCQR+Expansion shows superior performance in dense retrieval on the TopiOCQA, attaining the best MRR (38.5), NDCG (37.6), R@10 (58.4), and R@100 (75.0). The reported results show significant improvements with the $t$-test at $p<0.05$ overall compared to baselines in all the settings on QReCC and TopiOCQA (except AdaCQR+Expansion with ANCE on TopiOCQA).
@@
4.2 Zero-shot Results
To access the generalization performance of AdaCQR, we conduct the experiment on TREC CAsT datasets under a zero-shot setting, shown in Table 2. Among all non-expansion methods, AdaCQR achieves the best performance across all metrics on the CAsT-19 and the highest MRR scores on the CAsT-20 and CAsT-21. Among methods with expansion, AdaCQR still performs best on the CAsT-19 dataset and demonstrates comparable results to the LLM4CS-RAR approach on the CAsT-20 and CAsT-21. The strongest competitor in expansion setting, LLM4CS-RAR, leverages advanced proprietary LLM and candidates queries aggregation to enhance its expansion capability, so it is reasonable that AdaCQR falls slightly short. Achieving comparable performance further demonstrates the strength of our method. These results provide solid evidence of the generalization capabilities of our approach.
@@ 4.3 Ablation Study
Type Abaltion Variants MRR R@10
Sparse AdaCQR (Ours) 52.4 70.9
@@
Dense AdaCQR (Ours) 45.1 66.3
@@
Table 3: Ablation study for alignment and ranking of AdaCQR on QReCC dataset.
Type Query Reform. MRR R@10
Sparse Human Rewrite 39.8 62.7
LLM Rewrite $Q^{\star}$ 45.4 65.5
AdaCQR
   Trained on #HR 44.9 63.7
   Trained on $Q^{\star}$ 52.4 70.9
Dense Human Rewrite 38.4 58.6
LLM Rewrite $Q^{\star}$ 40.1 60.2
AdaCQR
   Trained on #HR 41.0 60.5
   Trained on $Q^{\star}$ 45.1 66.3
@@
Table 4: Ablation study for reformulation labels on QReCC dataset. The superior LLM reformulation $Q^{\star}$ is obtained by in-context learning (detailed in §3.4).
To investigate the impact of each component on the performance of AdaCQR, we conducted ablation experiments focusing on the below modules in Table 3 and Table 4.
Alignment
The training Stage of AdaCQR incorporates a contrastive loss to align the retrievers. To assess the influence of contrastive loss, we executed a single-stage training process without alignment. The removal of contrastive loss results in the most notable decline in performance, with a 9.1% decrease in MRR metric.
Ranking
We introduce a fusion metric to evaluate query performance across semantic and term perspectives. To determine the effect of the fusion metric, we substitute it with the metric used in previous work Jang et al. (2023), which ranks candidates solely based on the cosine similarity between the candidate query and the gold passage. The performance degradation confirms the effectiveness of using signals from the ranking of both types of retrievers. To further investigate the effectiveness of considering both perspectives in the fusion metric, we separately remove sparse ranking $r_{s}$ and dense ranking $r_{d}$ within it for analysis. Removing any rank degrades performance for both retrievers, more significantly for the corresponding retriever. This confirms the rationale behind considering both perspectives simultaneously.
Reformulation Labels
The performance of directly using human label and $Q^{\star}$ as the queries is reported in Table 4. It is worth noting that superior labels $Q^{\star}$ outperform human labels and have strong performance both in sparse and dense retrievals, which validates the effectiveness of the proposed fusion metric and the annotation method. In the Training Stage 1 of AdaCQR, we use $Q^{\star}$ as the golden labels. We also experiment by replacing it with human labels, and the resulting performance drop further validates the effectiveness of $Q^{\star}$. Additionally, the final results of AdaCQR after both Training Stage 1 and Stage 2 show significant improvement over the original queries (whether $Q^{\star}$ or human labels), demonstrating the effectiveness of the Stage 2 alignment.
4.4 Performance against CDR Methods
Conversational Dense Retrieval (CDR) is an orthogonal method for conversational query reformulation in conversational search, which trains dense retrievers to improve the representation of the current query and historical context. Although not directly comparable, we still present a performance comparison of the AdaCQR and CDR methods on the QReCC, TopiOCQA, and CAsT datasets, as shown in Table 5.
We compare our approach against four baseline models: Conv-ANCE Xiong et al. (2020), InstructoR-ANCE Jin et al. (2023), Conv-SPLADE Formal et al. (2021), and LeCoRE Mao et al. (2023c). Conv-ANCE trains a dense retriever by enhancing session representation using a conventional contrastive ranking loss. InstructoR-ANCE leverages large language models (LLMs) to predict the relevance score between sessions and passages, followed by retriever training. Conv-SPLADE fine-tunes a strong lexical-based retriever on conversational search data, utilizing ranking loss. LeCoRE extends the SPLADE model with multi-level denoising techniques to enhance lexical session representation. AdaCQR achieves the best average performance across four datasets, demonstrating superiority in both effectiveness and generalizability. Our approach employs the off-the-shelf ANCE retriever, and our rewrite method is orthogonal to the aforementioned retrievers, leaving the exploration of combining these methods for achieving higher performance in future research. Moreover, while CDR methods focus on dense representations, there are scenarios where the sparse representation of BM25 demonstrates a retrieval advantage. Our approach AdaCQR takes both types of retrievers into account and enhances performance through a rewrite strategy.
Method TopiOCQA QReCC CAsT-20 CAsT-21 Avg.
Conv-ANCE Xiong et al. (2020) 22.9 47.1 42.2 52.3 41.1
InstructoR-ANCE Jin et al. (2023) 25.3 43.5 43.7 53.0 41.4
Conv-SPLADE Formal et al. (2021) 30.7 50.0 36.9 47.9 41.4
LeCoRE Mao et al. (2023c) 32.0 51.1 37.7 50.8 42.9
AdaCQR(Ours) 32.8 45.1 51.4 58.3 46.9
AdaCQR+Expansion(Ours) 38.5 45.8 56.6 64.2 51.3
Table 5: MRR performance comparison of the AdaCQR and CDR methods.
@@ 5.1 Analysis of the Aligned Query
Figure 3: Analysis of the aligned reformulation query across epochs in Stage 2 training, focusing on the term overlap (DICE coefficient) and semantic similarity with the gold passage (cosine similarity).
To evaluate the effectiveness of the aligned reformulation queries, we analyse the reformulation queries across the first 5 epochs during Stage 2 training in Figure 3. We conduct analyses focusing on the average term overlap and semantic similarity between the queries and the gold passages. The DICE Coefficient Dice (1945) is utilized to assess term overlap, while cosine similarity is employed to measure semantic similarity. This analysis indicates that both term overlap and semantic similarity between the reformulated queries and the gold passages exhibit an increasing trend with each epoch in Stage 2, demonstrating the effectiveness of our method in considering both perspectives.
@@ 6 Conclusion
In this paper, to achieve alignment between the reformulation model and both term and semantic retrieval systems, AdaCQR is proposed to enhance the generalizability of information-seeking queries through a two-stage training strategy. By leveraging a fusion metric that assesses the generalization performance across different retrieval systems, we can effectively obtain superior labels for the generation and gather a diverse set of ordered candidate queries for reference-free evaluation. Extensive experiments on five datasets demonstrate the superiority of AdaCQR, achieving performance comparable to methods using the fine-tuned LLaMA2-7B and proprietary LLM.
@@
During the AdaCQR training process, we leverage ChatGPT for superior reformulation label annotation, and our annotation prompt requires training a basic model, which incurs additional costs and training expenses. Furthermore, due to budget constraints, we do not use more powerful LLMs, such as GPT-4, to obtain reformulation labels, although it is obvious that employing a more powerful LLM would yield better reformulation labels.
Although no further costs are introduced during reformulation model inference, aligning AdaCQR with retrievers introduces additional training time. Furthermore, generating the ordered candidate set for alignment demands extra retrieval time and increased storage capacity.
Acknowledgments
The authors would like to thank the anonymous reviewers for their insightful comments. This work is funded by the National Natural Science Foundation of China (Grant No.62176053). This work is supported by the Big Data Computing Center of Southeast University.